Karlottalotta:

Max observes the zoo and the library from a helicopter flying at a height of 300 times square root of 3 feet above the ground, as shown below: A helicopter is flying at a height of 300 multiplied by square root of 3 feet above the ground. A zoo and a library are on the What is the distance between the zoo and the library? (1 point) 300 feet 600 feet 900 feet 100 feet Im pretty sure its either b or c

jhonyy9:

as shown - please post this image h = 300 *sqrt3

jhonyy9:

@Laylalyssa

Laylalyssa:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @jhonyy9 @Laylalyssa \(\color{#0cbb34}{\text{End of Quote}}\) idk how to do this one either-

jhonyy9:

ok but the image is missed from there - yes ?

Laylalyssa:

@Karlottalotta is an image available for this question ? if so post it now...please <3

YECR:

DId u solve it or do you still need help?..

supie:

Asked & Answered here.

AZ:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @jhonyy9 ok but the image is missed from there - yes ? \(\color{#0cbb34}{\text{End of Quote}}\) You were right. From the post that @supie linked, we get the remaining information that helps us draw the image: A helicopter is flying at a height of 300 multiplied by square root of 3 feet above the ground. A zoo and a library are on the ground on the same side of the helicopter. The angle made by the line joining the helicopter and the zoo with the ground is 60 degrees. The angle made by the line joining the helicopter and the library with the ground is 30 degrees. What is the distance between the zoo and the library?

AZ:

So @Karlottalotta do you know some basic trigonometry? \(\sin(x) = \dfrac{\text{opposite side}}{\text{hypotenuse}}\) \(\cos(x) = \dfrac{\text{adjacent side}}{\text{hypotenuse}}\) \(\tan(x) = \dfrac{\text{opposite side}}{\text{adjacent side}}\)

1 attachment